A 2.59 ? resistor, a 5.83 ? resistor, and a 8.76 ? resistor are connected in series with a 12.0 V battery. What is the equivalent resistance?

What is the current in each resistor?
Answer #1

a)
R equivalent = R1 +R2 +R3
                    = 2.59 + 5.83+8.76
                    =17.18 ohm

b)
current in each resistor = V/ Requivalent
                       = 12 / 17.18
                           =0.7 A